# Phase-Field simulations of dynamic wetting by the Volume-Of-Fluid method

### v1

@@ 4 The toy model: relating mass flux to curvature
$\rho_{l}\left(\dfrac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\right)=-\dfrac{1}{\operatorname{Re}}\nabla P^{*}+\dfrac{1}{\operatorname{Re}}\nabla^{2}\mathbf{u}-\dfrac{1}{\operatorname{Re}\>\operatorname{Cn}\>\operatorname{Ca}}C\nabla\phi,$ (4.1)
$\dfrac{\partial C}{\partial t}+\mathbf{u}\cdot\nabla C=\dfrac{1}{\operatorname{Pe}}\nabla^{2}\phi,$ (4.2)
@@
$\phi=-\operatorname{Cn}^{2}\nabla^{2}C+\Psi^{\prime}\left(C\right),$ (4.3)
@@
$\Psi^{\prime}\left(C\right)=(C^{2}-1)C.$ (4.4)
@@
$\dfrac{3}{2\sqrt{2}}\operatorname{Cn}\nabla C\cdot n=\cos(\theta_{\mathrm{e}})g^{\prime}(C).$ (4.5)
@@
Assume a length scale $\ell$, a viscous balance in the momentum equation, and that the velocity is properly scaled (dimensionless $|\mathbf{u}|\sim 1)$. Estimates of the relevant terms in (4.1)- (4.2) then give
$\dfrac{1}{\operatorname{Re}}\>\dfrac{1}{\ell^{2}}\quad\sim\quad\dfrac{1}{\operatorname{Ca}\operatorname{Cn}\operatorname{Re}}\>\dfrac{\phi}{\ell},$ (4.6)
@@
$\dfrac{1}{\ell}\quad\sim\quad\frac{1}{Pe}\>\frac{\phi}{\ell^{2}}.$ (4.7)
@@
$\ell\quad\sim\quad\sqrt{\dfrac{\operatorname{Ca}\operatorname{Cn}}{\operatorname{Pe}}}=\dfrac{\sqrt{\mu_{l}M}}{L},$ (4.8)
which agrees with the scaling mentioned by several authors previously (Jacqmin (2000); Qian et al. (2006); Yue et al. (2010)). Briant & Yeomans (2004) has given a somewhat different expression $(\mu_{l}M\epsilon^{2})^{1/4}/L$. This however involves the interface thickness in the estimates of the concentration gradients, while the length scale in (4.8) characterizes the length scale for diffusion in the contact line region, away from the interface.
We now derive an expression for the chemical potential on a locally curved interface. Assuming a planar geometry and a circular interface of radius $R$, the chemical potential as defined in (4.3) can be written as
$\phi=-\operatorname{Cn}^{2}\left(\dfrac{\partial^{2}C}{\partial r^{2}}+\dfrac{1}{r}\dfrac{\partial C}{\partial r}\right)+\Psi^{\prime}\left(C\right).$ (4.9)
$C$ is expected to approach $+1$ inside the droplet, and $-1$ outside. Since the dimensionless interface width is $\operatorname{Cn}\ll 1$, the solution to this equation has a boundary layer character, and we introduce a new radial coordinate $s$ such that $r=R+\operatorname{Cn}\cdot s$ and expand (4.9) at zero and first order
$C=C_{0}+\operatorname{Cn}\cdot C_{1}+O(\operatorname{Cn}^{2}),$ (4.10)
$\phi=\phi_{0}+\operatorname{Cn}\cdot\phi_{1}+O(\operatorname{Cn}^{2}).$ (4.11)
@@
$\phi_{0}=-\dfrac{\partial^{2}C_{0}}{\partial s^{2}}+\Psi^{\prime}\left(C_{0}\right),$ (4.12)
@@
$\phi_{1}=-\left(\dfrac{\partial^{2}C_{1}}{\partial s^{2}}-\dfrac{1}{R}\dfrac{\partial C_{0}}{\partial s}\right)+C_{1}\Psi^{\prime\prime}\left(C_{0}\right).$ (4.13)
The zeroth order (4.12) gives the equilibrium solution as
$C_{0}=-\tanh(s/\sqrt{2}),\text{ with }\phi_{0}=0.$ (4.14)
Multiplying (4.13) by $\partial C_{0}/\partial s$ and partially integrating, a solvability condition is obtained for the first-order value of the chemical potential
@@
where the subscript denotes derivative with respect to $s$. The chemical potential is continuous at the interface, so the first order chemical potential $\phi_{1}$ is a constant in (4.13) and (4.15). By partially integrating the first term on the right hand side that term is seen to be zero. Using that
@@
$\phi_{i}=\operatorname{Cn}\phi_{1}=\dfrac{1}{2}\dfrac{2\sqrt{2}}{3}\dfrac{\operatorname{Cn}}{R}.$ (4.17)
Turning now to (4.2), which expresses the mass balance of the concentration $C$, the solution to this equation varies on two different length-scales: one is the interface width and the other a larger length scale characterizing the region around the contact line where mass diffusion restores the concentration to its equilibrium value, i.e. the Yue-Zhou-Feng length-scale $\ell$ in (4.8). In order to establish a quantitative model, we now rewrite the scaling of (4.2) in (4.7) as an equality
$b\>\dfrac{2}{\delta}\quad=\quad\frac{1}{\operatorname{Pe}}\>\frac{\phi_{i}}{\delta^{2}}.$ (4.18)
Here we introduced a non-dimensional factor $b$ of order unity, which is to account for the shape of the velocity profile at the contact line. We note that the full equation (4.2), when re-scaled according to (4.8), will be a non-dimensional equation, independent of model parameters. As $b$ is determined from a solution of this re-scaled system, we conclude that $b$ is independent of the model parameters; with the exception of the equilibrium contact angle $\theta_{\mathrm{e}}$. For partially wetting systems where $\theta_{\mathrm{e}}$ is not close to 0 or $180\degree$ we expect this dependence to be weak.
We now consider the Navier-Stokes equations (4.1). Assuming a balance between the viscous term and the capillary term, we estimate the relevant terms in the Navier-Stokes equation as
$\dfrac{1}{\operatorname{Re}}\,\dfrac{\tau}{\delta}\quad=\quad\dfrac{1}{\operatorname{Ca}\operatorname{Cn}\operatorname{Re}}\,\phi_{i}\,\dfrac{2}{\delta}.$ (4.19)
@@
Solving now for $R$ and $\delta$ from (4.17)- (4.19) yields the local curvature as a function of the flow parameters
$R=\dfrac{1}{2}\dfrac{2\sqrt{2}}{3}\dfrac{1}{\sqrt{ab}}\sqrt{\dfrac{\operatorname{Cn}}{\operatorname{Pe}\operatorname{Ca}}}.$ (4.20)
$\delta=\dfrac{1}{2}\sqrt{\dfrac{a}{b}}\sqrt{\dfrac{\operatorname{Ca}\operatorname{Cn}}{\operatorname{Pe}}}.$ (4.21)
Throughout the treatment above, the viscosity is taken as the higher liquid viscosity, in the case of unequal viscosities in the two regions, and as noted above, the dependence of our results on the viscosity ratio is weak. This may seem to be in contrast with the observation of Yue et al. (2010), Figures 9 and 11, who note that the distance from the wall to the stagnation point near the contact line (the boundary layer thickness) should be based on the geometric mean viscosity $\mu_{e}=\sqrt{\mu_{l}\,\mu_{a}}$. Here however the interpretation of the boundary layer thickness is not the distance to the stagnation point per se, but rather estimates the width of the highly curved region. One explanation could be that, as is clear from flow fields in the phase field simulations, the stagnation point is displaced into the low viscosity region when the viscosities differ. For the analysis here it is the action of the viscous stresses and the shape of the velocity profile in the contact line region that is important, and not the location of the stagnation point as such.
The local radius of curvature $R$ at the contact line allows us to estimate how the local angle $\theta_{\mathrm{a}}$ varies as function of the distance to the wall. Assuming a locally circular interface of radius $R$ according to (4.20), some geometrical considerations show that the angle $\theta_{\mathrm{a}}$, taken at a distance $\delta$ from the wall according to (4.21), is given by the following relation:
$\cos(\theta_{\mathrm{a}})=\cos(\theta_{\mathrm{e}})+\dfrac{\delta}{R}=\cos(\theta_{\mathrm{e}})+\dfrac{3}{2\sqrt{2}}\operatorname{Ca}a.$ (4.22)
In what follows we will use (4.22) to calculate the apparent contact angle $\theta_{\mathrm{a}}$ from the local capillary number $\operatorname{Ca}$, and apply this at a fictitious boundary displaced the distance $\delta$ into the flow domain, as indicated in Figure 1.
5 Results
5.1 Sheared drop
@@
Figure 2: Correlation between the radius of curvature $R$ and (4.20). (a) Each simulation corresponds to a unique set of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers while maintaining a constant capillary number ($\operatorname{Ca}$). The geometrical factor $a$ remains fixed at 2.5, and a fitted line demonstrates a clear one-to-one relationship between these two parameters. (b) Steady-state solutions of Phase-Field (PF) simulations for cases highlighted in red in the left panel.
In the first test case, we study the steady shape of a nanoscopic droplet confined between two moving walls similarly to Lācis et al. (2020, 2022). Initially, we use the Phase-Field (PF) model to solve the system in order to validate the correlation between the radius of curvature and the flow parameters, as expressed in (4.20). We choose a constant capillary number $\operatorname{Ca}=0.0212$, which is below the critical capillary number, while systematically varying the Peclet and Cahn numbers. The geometrical factor $a$ is kept fixed at a value of 2.5. For each simulation run, we compute the radius of curvature $R$ once the system attains a steady state, using the mean curvature at the contact line. The results, illustrated in Figure 2, demonstrate a remarkable agreement between $R$ and (4.20), establishing a clear one-to-one relationship between these two parameters. This preliminary analysis reinforces our choice of the geometrical factor $a$. In Section 5.2, we will further substantiate the appropriateness of this choice in the context of a spreading droplet.
We now test our toy model on this setup. Using the relations described in Section 4, we compare full PF simulations – in the sense that the full domain is solved – with VOF simulations using the toy model approximation. In the toy model, the bottom boundary is now a computational boundary that is located a distance $+\delta$ away from the wall. The domain comprised between the wall $y=0$ and the computational boundary $y=\delta$ is not simulated. There, we apply the dynamic angle relation (4.22), where the apparent angle $\theta_{\mathrm{a}}$ is now a dynamic contact angle $\theta_{d}$
$\theta_{d}=\cos^{\shortminus 1}\left(\cos\>\theta_{\mathrm{e}}+\dfrac{3}{2\sqrt{2}}\operatorname{Ca}a\right),$ (5.1)
@@
$u_{x}|_{y=\delta}-\delta\dfrac{\partial u_{x}}{\partial y}|_{y=\delta}=\pm U_{w},$ (5.2)
with $U_{w}$ the wall velocity. Figure 3 shows the resulting steady state interfaces for the left side of a sheared droplet, for a range of $\operatorname{Pe}$ and $\operatorname{Cn}$. The full PF simulations are compared to the corresponding VOF results, using equations (5.1) and (5.2) as boundary conditions, with good agreement. A more discerning test is shown in Figure 4 where the local angle that the interface makes with the horizontal is shown as a function of vertical position, for the same six cases as in Figure 3. The details of the profile shapes are near identical in all cases.
Figure 3: Comparison between the Phase-Field (PF) left interface (black points) and the Volume-of-Fluid (VOF) interface (red line) for six distinct sets of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers. Panels (a) to (f) correspond to the following parameter sets: ($\operatorname{Pe}=30,\>\operatorname{Cn}=0.02$), ($\operatorname{Pe}=30,\>\operatorname{Cn}=0.04$), ($\operatorname{Pe}=100,\>\operatorname{Cn}=0.02$), ($\operatorname{Pe}=160,\>\operatorname{Cn}=0.04$), ($\operatorname{Pe}=300,\>\operatorname{Cn}=0.02$), and ($\operatorname{Pe}=300,\>\operatorname{Cn}=0.04$).
(a) $\operatorname{Pe}=30$
$\operatorname{Cn}=0.02$
(b) $\operatorname{Pe}=30$
$\operatorname{Cn}=0.04$
(c) $\operatorname{Pe}=100$
$\operatorname{Cn}=0.02$
(d) $\operatorname{Pe}=160$
$\operatorname{Cn}=0.04$
(e) $\operatorname{Pe}=300$
$\operatorname{Cn}=0.02$
(f) $\operatorname{Pe}=300$
$\operatorname{Cn}=0.04$
Figure 4: Comparison between the Phase-Field (PF) angle along the left interface (black points) and the Volume-of-Fluid (VOF) angle (red line) for six distinct sets of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers.
5.2 Spreading drop

### v2

@@ IV The toy model: relating mass flux to curvature
$\rho_{l}\left(\dfrac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\right)=-\dfrac{1}{\operatorname{Re}}\nabla P^{*}+\dfrac{1}{\operatorname{Re}}\nabla^{2}\mathbf{u}-\dfrac{1}{\operatorname{Re}\>\operatorname{Cn}\>\operatorname{Ca}}C\nabla\phi,$ (22)
$\dfrac{\partial C}{\partial t}+\mathbf{u}\cdot\nabla C=\dfrac{1}{\operatorname{Pe}}\nabla^{2}\phi,$ (23)
@@
$\phi=-\operatorname{Cn}^{2}\nabla^{2}C+\Psi^{\prime}\left(C\right),$ (24)
@@
$\Psi^{\prime}\left(C\right)=(C^{2}-1)C.$ (25)
@@
$\dfrac{3}{2\sqrt{2}}\operatorname{Cn}\nabla C\cdot n=\cos(\theta_{\mathrm{e}})g^{\prime}(C).$ (26)
@@
Assume a length scale $\ell$, a viscous balance in the momentum equation, and that the velocity is properly scaled (dimensionless $|\mathbf{u}|\sim 1)$. Estimates of the relevant terms in (22)- (23) then give
$\dfrac{1}{\operatorname{Re}}\>\dfrac{1}{\ell^{2}}\quad\sim\quad\dfrac{1}{\operatorname{Ca}\operatorname{Cn}\operatorname{Re}}\>\dfrac{\phi}{\ell},$ (27)
@@
$\dfrac{1}{\ell}\quad\sim\quad\frac{1}{Pe}\>\frac{\phi}{\ell^{2}}.$ (28)
@@
$\ell\quad\sim\quad\sqrt{\dfrac{\operatorname{Ca}\operatorname{Cn}}{\operatorname{Pe}}}=\dfrac{\sqrt{\mu_{l}M}}{L},$ (29)
which agrees with the scaling mentioned by several authors previously [32, 38, 18]. Briant et al. [39] have given a somewhat different expression $(\mu_{l}M\epsilon^{2})^{1/4}/L$. This however involves the interface thickness in the estimates of the concentration gradients, while the length scale in (29) characterizes the length scale for diffusion in the contact line region, away from the interface.
We now derive an expression for the chemical potential on a locally curved interface. Assuming a planar geometry and a circular interface of radius $R$, the chemical potential as defined in (24) can be written as
$\phi=-\operatorname{Cn}^{2}\left(\dfrac{\partial^{2}C}{\partial r^{2}}+\dfrac{1}{r}\dfrac{\partial C}{\partial r}\right)+\Psi^{\prime}\left(C\right).$ (30)
$C$ is expected to approach $+1$ inside the droplet, and $-1$ outside. Since the dimensionless interface width is $\operatorname{Cn}\ll 1$, the solution to this equation has a boundary layer character, and we introduce a new radial coordinate $s$ such that $r=R+\operatorname{Cn}\cdot s$ and expand (30) at zero and first order
$C=C_{0}+\operatorname{Cn}\cdot C_{1}+O(\operatorname{Cn}^{2}),$ (31)
$\phi=\phi_{0}+\operatorname{Cn}\cdot\phi_{1}+O(\operatorname{Cn}^{2}).$ (32)
@@
$\phi_{0}=-\dfrac{\partial^{2}C_{0}}{\partial s^{2}}+\Psi^{\prime}\left(C_{0}\right),$ (33)
@@
$\phi_{1}=-\left(\dfrac{\partial^{2}C_{1}}{\partial s^{2}}-\dfrac{1}{R}\dfrac{\partial C_{0}}{\partial s}\right)+C_{1}\Psi^{\prime\prime}\left(C_{0}\right).$ (34)
The zeroth order (33) gives the equilibrium solution as
$C_{0}=-\tanh(s/\sqrt{2}),\text{ with }\phi_{0}=0.$ (35)
Multiplying (34) by $\partial C_{0}/\partial s$ and partially integrating, a solvability condition is obtained for the first-order value of the chemical potential
@@
where the subscript denotes derivative with respect to $s$. The chemical potential is continuous at the interface, so the first order chemical potential $\phi_{1}$ is a constant in (34) and (36). By partially integrating the first term on the right hand side that term is seen to be zero. Using that
@@
$\phi_{i}=\operatorname{Cn}\phi_{1}=\dfrac{1}{2}\dfrac{2\sqrt{2}}{3}\dfrac{\operatorname{Cn}}{R}.$ (38)
Turning now to (23), which expresses the mass balance of the concentration $C$, the solution to this equation varies on two different length-scales: one is the interface width and the other a larger length scale characterizing the region around the contact line where mass diffusion restores the concentration to its equilibrium value, i.e. the Yue-Zhou-Feng length-scale $\ell$ in (29). In order to establish a quantitative model, we now rewrite the scaling of (23) in (28) as an equality
$b\>\dfrac{2}{\delta}\quad=\quad\frac{1}{\operatorname{Pe}}\>\frac{\phi_{i}}{\delta^{2}}.$ (39)
Here we introduced a non-dimensional factor $b$ of order unity, which is to account for the shape of the velocity profile at the contact line. We note that the full equation (23), when re-scaled according to (29), will be a non-dimensional equation, independent of model parameters. As $b$ is determined from a solution of this re-scaled system, we conclude that $b$ is independent of the model parameters; with the exception of the equilibrium contact angle $\theta_{\mathrm{e}}$. For partially wetting systems where $\theta_{\mathrm{e}}$ is not close to 0 or $180\degree$ we expect this dependence to be weak.
We now consider the Navier-Stokes equations (22). Assuming a balance between the viscous term and the capillary term, we estimate the relevant terms in the Navier-Stokes equation as
$\dfrac{1}{\operatorname{Re}}\,\dfrac{\tau}{\delta}\quad=\quad\dfrac{1}{\operatorname{Ca}\operatorname{Cn}\operatorname{Re}}\,\phi_{i}\,\dfrac{2}{\delta}.$ (40)
@@
Solving now for $R$ and $\delta$ from (38)- (40) yields the local curvature as a function of the flow parameters
$R=\dfrac{1}{2}\dfrac{2\sqrt{2}}{3}\dfrac{1}{\sqrt{ab}}\sqrt{\dfrac{\operatorname{Cn}}{\operatorname{Pe}\operatorname{Ca}}}.$ (41)
$\delta=\dfrac{1}{2}\sqrt{\dfrac{a}{b}}\sqrt{\dfrac{\operatorname{Ca}\operatorname{Cn}}{\operatorname{Pe}}}.$ (42)
Throughout the treatment above, the viscosity is taken as the higher liquid viscosity, in the case of unequal viscosities in the two regions, and as noted above, the dependence of our results on the viscosity ratio is weak. This may seem to be in contrast with the observation of Yue et al. [18], Figures 9 and 11, who note that the distance from the wall to the stagnation point near the contact line (the boundary layer thickness) should be based on the geometric mean viscosity $\mu_{e}=\sqrt{\mu_{l}\,\mu_{a}}$. Here however the interpretation of the boundary layer thickness is not the distance to the stagnation point per se, but rather estimates the width of the highly curved region. One explanation could be that, as is clear from flow fields in the phase field simulations, the stagnation point is displaced into the low viscosity region when the viscosities differ. For the analysis here it is the action of the viscous stresses and the shape of the velocity profile in the contact line region that is important, and not the location of the stagnation point as such.
The local radius of curvature $R$ at the contact line allows us to estimate how the local angle $\theta_{\mathrm{a}}$ varies as function of the distance to the wall. Assuming a locally circular interface of radius $R$ according to (41), some geometrical considerations show that the angle $\theta_{\mathrm{a}}$, taken at a distance $\delta$ from the wall according to (42), is given by the following relation:
$\cos(\theta_{\mathrm{a}})=\cos(\theta_{\mathrm{e}})+\dfrac{\delta}{R}=\cos(\theta_{\mathrm{e}})+\dfrac{3}{2\sqrt{2}}\operatorname{Ca}a.$ (43)
In what follows we will use (43) to calculate the apparent contact angle $\theta_{\mathrm{a}}$ from the local capillary number $\operatorname{Ca}$, and apply this at a fictitious boundary displaced the distance $\delta$ into the flow domain, as indicated in Figure 1.
V Results
V.1 Sheared drop
@@
Figure 2: Correlation between the radius of curvature $R$ and (41). (a) Each simulation corresponds to a unique set of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers while maintaining a constant capillary number ($\operatorname{Ca}$). The geometrical factor $a$ remains fixed at 2.5, and a fitted line demonstrates a clear one-to-one relationship between these two parameters. (b) Steady-state solutions of phase-field (PF) simulations for cases highlighted in red in the left panel.
In the first test case, we study the steady shape of a nanoscopic droplet confined between two moving walls similarly to Lacis et al. [19, 20]. Initially, we use the phase-field (PF) model to solve the system in order to validate the correlation between the radius of curvature and the flow parameters, as expressed in (41). We choose a constant capillary number $\operatorname{Ca}=0.0212$, which is below the critical capillary number, while systematically varying the Peclet and Cahn numbers. The geometrical factor $a$ is kept fixed at a value of 2.5. For each simulation run, we compute the radius of curvature $R$ once the system attains a steady state, using the mean curvature at the contact line. The results, illustrated in Figure 2, demonstrate a remarkable agreement between $R$ and (41), establishing a clear one-to-one relationship between these two parameters. This preliminary analysis reinforces our choice of the geometrical factor $a$. In Section V.2, we will further substantiate the appropriateness of this choice in the context of a spreading droplet.
We now test our toy model on this setup. Using the relations described in Section IV, we compare full PF simulations – in the sense that the full domain is solved – with VOF simulations using the toy model approximation. In the toy model, the bottom boundary is now a computational boundary that is located a distance $+\delta$ away from the wall. The domain comprised between the wall $y=0$ and the computational boundary $y=\delta$ is not simulated. There, we apply the dynamic angle relation (43), where the apparent angle $\theta_{\mathrm{a}}$ is now a dynamic contact angle $\theta_{d}$
$\theta_{d}=\cos^{\shortminus 1}\left(\cos\>\theta_{\mathrm{e}}+\dfrac{3}{2\sqrt{2}}\operatorname{Ca}a\right),$ (44)
@@
$u_{x}|_{y=\delta}-\delta\dfrac{\partial u_{x}}{\partial y}|_{y=\delta}=\pm U_{w},$ (45)
with $U_{w}$ the wall velocity. Figure 3 shows the resulting steady state interfaces for the left side of a sheared droplet, for a range of $\operatorname{Pe}$ and $\operatorname{Cn}$. The full PF simulations are compared to the corresponding VOF results, using equations (44) and (45) as boundary conditions, with good agreement. A more discerning test is shown in Figure 4 where the local angle that the interface makes with the horizontal is shown as a function of vertical position, for the same six cases as in Figure 3. The details of the profile shapes are near identical in all cases.
Figure 3: Comparison between the phase-field left interface (black points) and the volume-of-fluid interface (red line) for six distinct sets of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers. Panels (a) to (f) correspond to the following parameter sets: ($\operatorname{Pe}=30,\>\operatorname{Cn}=0.02$), ($\operatorname{Pe}=30,\>\operatorname{Cn}=0.04$), ($\operatorname{Pe}=100,\>\operatorname{Cn}=0.02$), ($\operatorname{Pe}=160,\>\operatorname{Cn}=0.04$), ($\operatorname{Pe}=300,\>\operatorname{Cn}=0.02$), and ($\operatorname{Pe}=300,\>\operatorname{Cn}=0.04$).
(a) $\operatorname{Pe}=30$ $\operatorname{Cn}=0.02$
(b) $\operatorname{Pe}=30$ $\operatorname{Cn}=0.04$
(c) $\operatorname{Pe}=100$ $\operatorname{Cn}=0.02$
(d) $\operatorname{Pe}=160$ $\operatorname{Cn}=0.04$
(e) $\operatorname{Pe}=300$ $\operatorname{Cn}=0.02$
(f) $\operatorname{Pe}=300$ $\operatorname{Cn}=0.04$
Figure 4: Comparison between the phase-field angle along the left interface (black points) and the volume-of-fluid angle (red line) for six distinct sets of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers.
V.2 Spreading drop